Translate into an equation and solve. Five times the difference between three times a number and four is ten. Find the number.
I get the basic concept, how to do it. But I keep getting different answers, in other words every time I do it it comes out a little different.
Thank You!

Answer by queenofit(120)

I think you would do it this way
5(3x-4)=10
15x-20=10
15x=20+10
15x=30
x=2